OpenStudy (anonymous):

Express this fraction in a+bi form: -5+i/1-6i Show work plz!

OpenStudy (anonymous):

\[\Large \frac{-5+i}{1-6i}\cdot \frac{1+6i}{1+6i}=\frac{(-5+i)(1+6i)}{1-36i^2}\] \[\Large \frac{(-5-30i+i+6i^2)}{1-36\cdot (-1)}=\frac{-11-29i}{37}\] \[\Large =\underbrace{\frac{-11}{37}}_{Re.}\;\;\;\;\;\;\;\;\underbrace{-\frac{29}{37}}_{Im.} \cdot i\] Instructions are given by Luis Rivera perfectly !